OpenStudy (anonymous):

Let S be the set of all integers that can be written as n2 + 1, where n is a nonzero integer. Which of the following integers is in S ?

OpenStudy (anonymous):

is it \[n ^{2}+1?\]

OpenStudy (anonymous):

@KNicksfan4life

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

answers?

OpenStudy (anonymous):

(A)16 (B)28 (C) 35 (D)39 (E) 50

OpenStudy (anonymous):

just notice 16=4^2 28=5^2+3 35=6^2-1 39=6^2+3 50=7^2+1

OpenStudy (anonymous):

what is ypur choice now?

OpenStudy (anonymous):

um what?

OpenStudy (anonymous):

can u choose the right answer using my hint?

OpenStudy (anonymous):

hell no

OpenStudy (anonymous):

can u just tell me what the answer is

OpenStudy (anonymous):

50

OpenStudy (anonymous):

thank u